OpenStudy (anonymous):

Use the three steps to solve the problem. Joe is 10 years older than James. In 8 years, twice Joe's age will equal three times James's age. How old is each now?

OpenStudy (anonymous):

Let "Joe" = Joe's current age, and "James" = James' current age. By the first constraint, (Equation 1) Joe = James + 10 By the second constraint, (Equation 2) 2 * (Joe + 8) = 3 * (James + 8) (Equation 3) 2 * Joe + 16 = 3 * James + 24 Subtracting 2 * (Equation 1) from (Equation 3), we get (Equation 4) 16 = James + 4 (Equation 5) James = 12 Substituting (Equation 5) back into (Equation 1), we see: (Equation 6) Joe = 12 + 10 = 22.